# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 7 A tree-boosting algorithm for parametric regression
Boosted tree algorithms have a long and successful history as robust solutions to various classification and regression tasks and were recently adapted to novel searches for generic resonant phenomena [54]. They consist of an additive sequence of weak learners, each providing a coarsely binned prediction based on a hierarchical phase-space partitioning that is fast to evaluate. The terminal nodes of each tree are associated with a predictive function that changes non-linearly across the phase-space boundaries of the terminal nodes. Technically, we extend standard tree-based regression algorithms, implemented, for example, in TMVA [55], with a more flexible terminal-node predictor that provides the DCR to arbitrary order in the expansion in ${\boldsymbol{\nu}}$. The summed prediction of the weak learners iteratively trained using a “boosting” algorithm, is smooth and arbitrarily expressive.
The weak learner’s simplicity results in relatively mild failure modes. For example, if a tree is trained with insufficient data, it cannot incorrectly extrapolate to phase space regions beyond the reach of the training data; it will result in the prediction of the highest bin the algorithm could find, respecting a regularization requirement, for example on the minimum number of events in each terminal node. Secondly, a tree algorithm with axis-linear node splits, such as produced by the “Classification and Regression Tree” (CART) [56] algorithm, does not interpolate step-wise features in the training data. Even if features take only discrete values, the optimal phase-space partitioning will be found in terms of selections at these discrete values. Features related to the multiplicity of reconstructed objects thus need no special treatment. Replacing the nominal training data with digitized values according to a binning choice does not present a conceptual challenge. Instead, such a modification could serve as validation when the sensitivity of an unbinned analysis is compared to a binned reference result.
The main perceived drawback of tree algorithms is related to the relative simplicity of the terminal nodes’ predictive function. We argue, however, that this is a failure of imagination. Nothing prevents us from associating more complex quantities with terminal nodes than the conventional class probabilities or regression values. Trees have been used to parametrically learn the polynomial SMEFT dependence [10, 9]. In the following sections, we develop a boosted tree algorithm where the weak learner’s terminal nodes are associated with parametrizations of systematic effects equivalent to the exponential in Eq. 56. Our goal is to provide an estimate $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})$ of the DCR
$\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})\simeq\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})\;\;\;}{{\textrm{d}}\sigma(\boldsymbol{x}|\textrm{SM})}$ (74)
that is fast to evaluate, parametric in ${\boldsymbol{\nu}}$, satisfies $\hat{S}(\boldsymbol{x}|{\mathbf{0}})=1$, and is continuous in $\boldsymbol{x}$ and ${\boldsymbol{\nu}}$.
A boosting algorithm iteratively fits the weak learners of an additive expansion to the pseudo-residuals of the preceding boosting iteration. During this training phase, each weak learner captures only a fraction of the total parameter dependence in each terminal node. The algorithm’s learning rate controls this fraction. The pseudo-residuals are synthetic data sets with the already learned dependence removed by a reweighting procedure. We start with tree boosting because it can be discussed independently from the weak learner’s details. In the subsequent sections, we detail the learning of the phase-space partitioning, construct the terminal node predictions, summarize the result in pseudo-code, and test it on a simple analytic toy example.
7.1 Tree-boosting of parametric regressors
It is instructive to discuss boosting for generic non-parametric estimators starting with the cross-entropy loss function $L_{\textrm{CE}}[\hat{f}]$ in Eq. 39 after the replacement in Eq. 47. We then have
$L[\hat{T}]$ $=\left\langle\textrm{Soft}^{+}(\hat{T}(\boldsymbol{x}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}}}+\left\langle\textrm{Soft}^{+}(-\hat{T}(\boldsymbol{x}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}},$ (75)
$T^{\ast}(\boldsymbol{x})=\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})\,},$ (76)
but instead of obtaining this result in a single fit, we chose a number $B$ of boosting iterations and corresponding learning rates $10<\eta^{(b)}<$ 1 for $b=1,\ldots,B$. We aim at an additive expansion of $\hat{T}(\boldsymbol{x})$ in terms of the weak learners $\hat{t}^{(b)}(\boldsymbol{x})$. To this end, we iterate the boosting relations
$t^{(b)}{}^{\ast}(\boldsymbol{x})$ $=\operatorname*{arg\,min}_{\hat{t}^{(b)}}L\left[\hat{t}^{(b)}{}(\boldsymbol{x})+\hat{T}^{(b-1)}(\boldsymbol{x})\right],$ (77)
$\hat{T}^{(b)}(\boldsymbol{x})$ $=\hat{T}^{(b-1)}(\boldsymbol{x})+\eta^{(b)}\hat{t}^{(b)}{}^{\ast}(\boldsymbol{x})$ (78)
a number of $B$ times, starting with the initial choice $\hat{T}^{(0)}(\boldsymbol{x})=0$. Equation 77 obtains the weak learner $\hat{t}^{(b)}(\boldsymbol{x})$ when the result of the preceding iteration $\hat{T}^{(b-1)}(\boldsymbol{x})$ is known. Equation 78 updates the additive model with a fraction $\eta^{(b)}$ of this weak learner’s prediction. After $B$ iterations, the boosted prediction $\hat{T}^{(B)}$ can be expressed as
$\hat{T}^{(B)}(\boldsymbol{x})$ $=\sum_{b=1}^{B}\eta^{(b)}t^{(b)}{}^{\ast}(\boldsymbol{x}).$ (79)
An important practicality for boosting learners that are fit to synthetic data sets follows from the minimum condition in Eq. 77. It implies that the minimum at iteration $b$ satisfies
$t^{(b)}{}^{\ast}(\boldsymbol{x})+\hat{T}^{(b-1)}(\boldsymbol{x})\simeq\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})}=\log\frac{\sigma({\boldsymbol{\nu}})}{\sigma({\mathbf{0}})}\frac{\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})}{\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}})}$ (80)
which we rearrange to
[…]
By reading this equation as an $\boldsymbol{x}$-dependent scaling of the joint-space integration measure ${\textrm{d}}\sigma(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})=\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}){\textrm{d}}\boldsymbol{x}{\textrm{d}}\boldsymbol{z}$ by the reciprocal of the estimate of the preceding boosting iteration, we find that $t^{(b)}{}^{\ast}$ can also be obtained if, instead of using the additive expansion, we replace $\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})\rightarrow\exp\left(-\hat{T}^{(b-1)}(\boldsymbol{x})\right)\,\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$. Because Eq. 78 provides the exponent $\hat{T}$ iteratively, we only have to multiply the cross-section by $\exp(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x}))$ when moving from iteration $b-1$ to iteration $b$. This way, the boosting equations read
$\sigma({\boldsymbol{\nu}})p^{(b)}(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$ $=\exp\left(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x})\right)\,\sigma({\boldsymbol{\nu}})p^{(b-1)}(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$ (82)
$t^{(b)}{}^{\ast}(\boldsymbol{x})$ $=\operatorname*{arg\,min}_{\hat{t}^{(b)}}L\left[\hat{t}^{(b)}{}(\boldsymbol{x})\right],$ (83)
$\hat{T}^{(b)}(\boldsymbol{x})$ $=\hat{T}^{(b-1)}(\boldsymbol{x})+\eta^{(b)}\hat{t}^{(b)}{}^{\ast}(\boldsymbol{x}),$ (84)
initialized by $\hat{T}^{(0)}(\boldsymbol{x})=t^{(0)}{}^{\ast}(\boldsymbol{x})=0$. The advantage of this formulation is that Eq. 83 is a standard loss function minimization without the additive model appearing in the argument as in Eq. 77. The update of the synthetic data set $\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}=\{w_{i}^{(b)},\boldsymbol{x}_{{\boldsymbol{\nu}},i},\boldsymbol{z}_{i}\}$, now also defined for each iteration $b$, follows from Eq. 82 as
$w_{i}^{(b)}=\exp\left(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x}_{i})\right)w^{(b-1)}_{i}.$ (85)
This prescription can be interpreted as a recursive weighting of the differential cross section of $\mathcal{D}_{{\boldsymbol{\nu}}}$ in the second term in Eq. 75 towards $\mathcal{D}_{{\mathbf{0}}}$ in the first term. The boosting algorithm removes the learned approximation from the training data as the regressor learns to approximate the DCR more accurately. It is customary to chose $\eta^{(b)}$ independently of $b$, and values between $10^{-3}$ and $3\cdot 10^{-1}$ for this universal learning rate have proven efficient.
The sample $\mathcal{D}_{{\mathbf{0}}}$ stays unchanged in the boosting procedure because we decided to write the $\boldsymbol{x}$-dependent scaling in Eq. 82 in the numerator. The choice of only reweighting the sample $\mathcal{D}_{\boldsymbol{\nu}}$ is a critical detail. It holds the key to a boosting algorithm that works for the fully parametric regressor, including the ${\boldsymbol{\nu}}$ dependence. We can construct the loss for a parametric tree-based algorithm from the general parametric loss function in Eq. 51, which is a sum of equally structured terms
$L$ $=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}L_{\textrm{CE}}[\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})]=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left(\left\langle\textrm{Soft}^{+}(\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}}}+\left\langle\textrm{Soft}^{+}(-\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}}\right).$ (86)
The synthetic data set in the first expectation value in each sum term is always $\mathcal{D}_{\mathbf{0}}$, irrespective of the value of ${\boldsymbol{\nu}}$. The synthetic data set in the second expectation is $\mathcal{D}_{{\boldsymbol{\nu}}}$ and is different for each $\mathcal{{\boldsymbol{\nu}}}\in\mathcal{V}$. It is this term whose synthetic data set changes during the boosting algorithm, and because there is one such set for each ${\boldsymbol{\nu}}\in\mathcal{V}$, the reweighting can be done simultaneously for each ${\boldsymbol{\nu}}$ in the sum over $\mathcal{V}$ in Eq. 86. Repeating the steps starting at Eq. 75 with a sufficiently expressive ${\boldsymbol{\nu}}$-dependent function $\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})$, it is straightforward to show that aside from the extra ${\boldsymbol{\nu}}$-dependence in the notation nothing else changes. Concretely, we only need to modify Eq. 79 to notate the ${\boldsymbol{\nu}}$-dependence in the weak learner $\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})$. The other steps follow analogously, and Eq. 85 generalizes to
$\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}=\left\{\exp(-\eta^{(b-1)}t^{(b-1)\ast}(\boldsymbol{x}_{i}|{\boldsymbol{\nu}})\,w^{(b-1)}_{i},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\right\}\;\;\textrm{for all}\;\;\{w^{(b-1)},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\}\in\mathcal{D}_{{\boldsymbol{\nu}}}^{(b-1)}\;\;\textrm{for all}\;\;{\boldsymbol{\nu}}\in\mathcal{V},$ (87)
which is the same as Eq. 85 except for that it is carried out simultaneously for each ${\boldsymbol{\nu}}\in\mathcal{V}$. This completes the boosting algorithm for generic weak learners, and we can proceed with constructing the tree-based implementation.
7.2 Learning the phase-space partitioning
$\mathcal{X}=\bigcup_{j\in\mathcal{J}}\Delta\boldsymbol{x}_{j}\quad\textrm{and}\quad\Delta\boldsymbol{x}_{j}\cap\Delta\boldsymbol{x}_{j^{\prime}}=\emptyset\;\;\longleftrightarrow\;\;j\neq j^{\prime}.$ (88)
$\mathbbm{1}_{j}(\boldsymbol{x})=\begin{cases}1&\textrm{if}\;\boldsymbol{x}\in\Delta\boldsymbol{x}_{j}\\ 0&\textrm{otherwise}\end{cases}$ (89)
$\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{j\in\mathcal{J}}\mathbbm{1}_{j}(\boldsymbol{x})\,\hat{t}_{j}({\boldsymbol{\nu}}).$ (90)
Because the $\boldsymbol{x}$-dependence is only in the index function, we can insert Eq. 90 into Eq. 86 and use Eq. 43 to carry out the event sums over the synthetic data sets. The result is
$L[\mathcal{J},\hat{t}_{j}]=\sum_{j\in\mathcal{J}}L_{j}[\hat{t}_{j}]=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,0}\,\textrm{Soft}^{+}(\hat{t}_{j}({\boldsymbol{\nu}}))+\sigma_{j,{\boldsymbol{\nu}}}\,\textrm{Soft}^{+}(-\hat{t}_{j}({\boldsymbol{\nu}}))\right].$ (91)
$\sigma_{j,0}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}_{\mathbf{0}}\cap\Delta\boldsymbol{x}_{j}}w_{i}\qquad\textrm{and}\qquad\sigma_{j,{\boldsymbol{\nu}}}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}_{\boldsymbol{\nu}}\cap\Delta\boldsymbol{x}_{j}}w_{i}$ (92)
Before we tackle these problems, it is instructive to develop an intuition for the loss function in Eq. 91. We assume an infinitely expressive $\hat{t}_{j}({\boldsymbol{\nu}})$ and functionally differentiate Eq. 91 to arrive at
$0=\frac{\delta L_{j}}{\delta\hat{t}_{j}}=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\frac{\sigma_{j,{\mathbf{0}}}}{1+\exp\left(-\hat{t}_{j}({\boldsymbol{\nu}})\right)}-\frac{\sigma_{j,{\boldsymbol{\nu}}}}{1+\exp\left(\hat{t}_{j}({\boldsymbol{\nu}})\right)}\right].$ (93)
This equation is satisfied exactly if
$\hat{t}_{j}({\boldsymbol{\nu}})=\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\quad\textrm{for all}\quad{\boldsymbol{\nu}}\in\mathcal{V},$ (94)
which we can always fulfill for sufficiently expressive $\hat{t}_{j}({\boldsymbol{\nu}})$; a perfect representation of the ${\boldsymbol{\nu}}$-dependence in each tree node $j\in\mathcal{J}$ reproduces the logarithm of the DCR for those values of ${\boldsymbol{\nu}}$ whose synthetic data sets we included in the loss function. The predictive function of a tree is finitely expressive. Hence, we will seek an approximation for Eq. 94 in the next section. But we can meanwhile use the result to shed light on the loss function Eq. 91. Formally eliminating $\hat{t}_{j}({\boldsymbol{\nu}})$ in favor of its predictions at the points ${\boldsymbol{\nu}}\in\mathcal{V}$, we express the loss solely in terms of the per-bin cross sections $\sigma_{j,{\mathbf{0}}}$ and $\sigma_{j,{\boldsymbol{\nu}}}$,
$L[\mathcal{J}]$ $=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,{\mathbf{0}}}\log\left(1+\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right)+\sigma_{j,{\boldsymbol{\nu}}}\log\left(1+\frac{\sigma_{j,{\mathbf{0}}}}{\sigma_{j,{\boldsymbol{\nu}}}}\right)\right].$ (95)
This equation would provide a loss function for finding the optimal phase space partitioning if we didn’t need to use finitely expressive $\hat{t}_{j}({\boldsymbol{\nu}})$. If we assume small ${\boldsymbol{\nu}}$ such that a Taylor expansion of $\sigma_{j,{\boldsymbol{\nu}}}$ around ${\boldsymbol{\nu}}={\mathbf{0}}$ is a good approximation, we get
$L[\mathcal{J}]$ $=-\frac{1}{4}\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\nu_{a}\nu_{b}I_{(ab),j}+\ldots$ (96)
$I_{(ab),j}=\frac{1}{\sigma_{j,{\mathbf{0}}}}\frac{\partial\sigma_{j,{\boldsymbol{\nu}}}}{\partial\nu_{a}}\frac{\partial\sigma_{j,{\boldsymbol{\nu}}}}{\partial\nu_{a}}\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}$ (97)
7.3 Terminal node predictions
$\hat{t}_{j}({\boldsymbol{\nu}})$ $=\nu_{a}\hat{\Delta}_{a,j}+\nu_{a}\nu_{b}\hat{\Delta}_{ab,j}+\nu_{a}\nu_{b}\nu_{c}\hat{\Delta}_{abc,j}+\cdots=\nu_{A}\hat{\Delta}_{A,j}$ (98)
$\hat{\Delta}_{j,a}\approx\frac{\partial}{\partial\nu_{a}}t_{j}({\boldsymbol{\nu}})\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}=\frac{\partial}{\partial\nu_{a}}\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})}\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}=s_{a}+\frac{\partial}{\partial\nu_{a}}\log\sigma({\boldsymbol{\nu}})\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}\quad\textrm{for}\quad\boldsymbol{x}\in\Delta\boldsymbol{x}_{j}.$ (99)
The last expression relates $\hat{\Delta}_{j,a}$ to the well-known score vector $s_{a}$, a sufficient statistic for small ${\boldsymbol{\nu}}$ and, therefore, an optimal observable. The log-derivative of the inclusive cross-section in the last term does not depend on the phase-space partitioning and, thus, is irrelevant to the optimization. The algorithm will, therefore, aim to reduce the expectation of the variance of the score in the training sample. This is, by definition, the negative value of the Fisher information matrix, consistent with the interpretation in the preceding section. Depending on the concrete problem and the desired accuracy, the higher-order terms in Eq. 99 can improve the parameterization for larger values of ${\boldsymbol{\nu}}$.
$\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{j\in\mathcal{J}}\mathbbm{1}_{j}(\boldsymbol{x})\left(\nu_{A}\hat{\Delta}_{A,j}\right).$ (100)
$\nu_{A}\hat{\Delta}_{A,j}$ $=\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\quad\textrm{for all}\quad{\boldsymbol{\nu}}\in\mathcal{V}$ (101)
are overdetermined if $|\mathcal{V}|>N_{\Delta}$. We note that Eq. 101 has the same form as Eq. 94 except for the finite expressivity on the l.h.s. We are content with an approximate solution of the per-node parametrization because boosting the weak learner will iteratively reduce the shortcomings either way. Any deficiency of a concrete weak learner will be reduced in the subsequent boosting iteration. For $|\mathcal{V}|<N_{\Delta}$, the training data cannot provide a unique estimate, and more data sets must be obtained.
The simplest approach for approximately solving Eq. 101 is by minimizing the mean-squared error separately for each node $j\in\mathcal{J}$,
$L_{j,\textrm{MSE}}[\hat{\Delta}]$ $=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left(\nu_{A}\hat{\Delta}_{A,j}-\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right)^{2}.$ (102)
It is solved by
$\hat{\Delta}_{A,j}=\left[\;\sum_{\nu\in\mathcal{V}}{\boldsymbol{\nu}}{\boldsymbol{\nu}}^{\textrm{T}}\right]_{AB}^{-1}\left[\sum_{\;{\boldsymbol{\nu}}\in\mathcal{V}}{\boldsymbol{\nu}}\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right]_{B}.$ (103)
The matrix
$V_{AB}=\left[\sum_{\nu\in\mathcal{V}}{\boldsymbol{\nu}}{\boldsymbol{\nu}}^{\textrm{T}}\right]_{AB},$ (104)
It is instructive to check that the weak learner appropriately responds to training data that is perfectly, not just approximately, consistent with the polynomial ansatz. If we take constants $\delta_{A}$ and consider a model that predicts $\sigma_{j,{\boldsymbol{\nu}}}=\exp(\nu_{A}\delta_{A})\sigma_{j,{\mathbf{0}}}$ in a given region, we can insert into Eq. 103 and find $\hat{\Delta}_{A}=\delta_{A}$, confirming that the algorithm learns the exact solution if it has the chance.
To complete the construction of the weak learner, we insert the ansatz Eq. 100 into Eq. 91 and get
$L[\mathcal{J}]$ $=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,{\mathbf{0}}}\,\textrm{Soft}^{+}\left(\nu_{A}\hat{\Delta}_{j,A}\right)+\sigma_{j,{\boldsymbol{\nu}}}\,\textrm{Soft}^{+}\left(-\nu_{A}\hat{\Delta}_{j,A}\right)\right],$ (105)
where $\hat{\Delta}_{A,j}$ are obtained from Eq. 103 and $\sigma_{j,0}$ and $\sigma_{j,{\boldsymbol{\nu}}}$ from Eq. 92. The data samples $\mathcal{D}_{\mathbf{0}}$, used for the prediction of $\sigma_{j,{\mathbf{0}}}$ in the first term, can either taken to be the same or statistically independent samples. This loss function is amenable to standard tree algorithms, for example, the CART algorithm or the “Tree Alternate Optimization” (TAO) [57, 58, 59, 60] algorithm, both providing tree structures with a hierarchical selection using the features $\boldsymbol{x}$ and that satisfy the requirements in Eq. 88. These algorithms proceed by recursively splitting the training data along either axis-aligned (for CART) or linear combinations of the input features (for TAO), reducing the loss at each iteration. The maximum recursion depth and the minimum number of events in each node are hyperparameters that regularize the fit. If no more splits can be performed, the terminal selections (nodes) represent a phase space partitioning $\mathcal{J}$ of the form in Eq. 88 and the quantities $\hat{\Delta}_{j}$ can be computed from Eq. 103. The tree then estimates the (log-)DCR as in Eq. 100. As a function of $\boldsymbol{x}$, the prediction of a single tree changes discontinuously if $\boldsymbol{x}$ traverses a boundary between nodes, and the possibly poor approximation close to the boundaries weakens the learner. Utilizing boosting, i.e., using a sequence of trees in Eq. 77 and Eq. 78, we recover the smooth behavior in $\boldsymbol{x}$ of an arbitrarily expressive regressor. Because each tree in the boosted result in Eq. 79 is parametric in ${\boldsymbol{\nu}}$, so is the final parametric regression tree.
7.4 Algorithm summary
Algorithm 1 “Parametric regression tree” for learning of systematic uncertainties
@@ 7.4 Algorithm summary
We can combine the steps in Secs. 7.17.3 and summarize the parametric regression tree algorithm. It is an iterative fit of a tree-based weak learner with the loss in Eq. 105 to the residuals of the preceding boosting iteration whose predictions are obtained from Eqs. 8285. Concretely, we start with training data $\mathcal{D}_{0}$ at a reference point and several synthetic data sets associated with model parameters ${\boldsymbol{\nu}}\in\mathcal{V}$. We must have enough data so $V_{AB}$ has full rank. We fit a weak learner using the CART algorithm. At each recursion step, the CART algorithm recursively divides the feature space by greedily selecting the dimension and cut value combination that minimizes the loss function. Overfitting is mitigated by enforcing a maximum tree depth and a minimum number of events in each terminal node. We construct new synthetic data from the weak learner’s prediction using Eq. 85 to replace $\mathcal{D}_{\boldsymbol{\nu}}$. This reweighting procedure brings the samples $\mathcal{D}_{\boldsymbol{\nu}}$ closer to $\mathcal{D}_{0}$ by an amount controlled by the learning rate $\eta$. We iterate the whole procedure $B$ times and obtain the final result from Eq. 79 as
$\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})=\log\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{b=1}^{B}\eta^{(b)}\sum_{j\in\mathcal{J}^{(b)}}\mathbbm{1}_{j}(\boldsymbol{x})\nu_{A}\hat{\Delta}^{(b)}_{A,j},$ (106)
7.5 An analytic toy example
To illustrate the parametric regression tree, we consider an arbitrarily chosen one-dimensional two-parameter model

### v3

@@ 5.4 The Boosted Parametric Tree algorithm
Boosted tree algorithms have a strong track record in classification and regression tasks and were recently applied to novel searches for resonant phenomena [52]. They use an additive sequence of weak learners, each generating a coarsely binned prediction based on hierarchical phase-space partitioning, which is computationally efficient. Each tree’s terminal nodes are linked to a predictive function that varies non-linearly across the phase-space boundaries of these nodes. Here, we extend standard tree-based regression algorithms, such as those in TMVA [53], by introducing a more flexible terminal-node predictor that provides the DCR to arbitrary order in the expansion in ${\boldsymbol{\nu}}$. The summed prediction from these weak learners, trained iteratively through boosting, is both smooth and arbitrarily expressive.
The simplicity of the weak learner leads to relatively mild failure modes. For instance, if a tree is trained with insufficient data, it cannot extrapolate incorrectly to phase-space regions beyond the training data; it will simply predict the value of the highest populated bin it finds, respecting regularization requirements, such as a minimum number of events per terminal node.
Additionally, a tree algorithm with axis-linear node splits, like the “Classification and Regression Tree” (CART) [54] algorithm, does not interpolate stepwise features in the training data. If features take only discrete values, the algorithm will partition phase space based on selections at these discrete values. Therefore, features related to object multiplicity need no special handling. Replacing nominal training data with digitized values according to a chosen binning can serve as validation, allowing sensitivity comparisons between unbinned and binned reference results.
The trees’ terminal nodes can be associated with more complex quantities than simple class probabilities or regression values, as shown in applications learning polynomial SMEFT dependence [10, 9]. We exploit this flexibility for developing a boosted tree algorithm where terminal nodes of the weak learner are linked to parametrizations of systematic effects of the training data in the nodes. With the tools from Sec. 5.2, the BPT provides a tree-based estimate $\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})$ of the logarithm of the DCR in the polynomial expansion
$\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})=\nu_{A}\hat{\Delta}_{A}(\boldsymbol{x})\simeq\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})\;\;\;}{{\textrm{d}}\sigma(\boldsymbol{x}|\textrm{SM})}$ (56)
so that $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\exp(\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))$. The free parameters in $\hat{\Delta}_{A}(\boldsymbol{x})$ are trained with the CE loss function in Eq. 52 by an iterative boosting algorithm. It fits the weak learners of an additive expansion, one at a time, to the pseudo-residuals of the preceding boosting iteration. The complete construction is discussed in Appendix C. Here, we describe the resulting algorithm.
During training, each weak learner captures only part of the total parameter dependence in each terminal node, with this fraction controlled by the algorithm’s learning rate. We set a number $B$ of boosting iterations and choose learning rates $10<\eta^{(b)}<$ 1 for $b=1,\ldots,B$ to form an additive expansion of $\hat{T}(\boldsymbol{x})$ in terms of the weak learners. The $\eta^{(b)}$ can be chosen constant, and values between $10^{-3}$ and $3\cdot 10^{-1}$ for this universal learning rate have proven efficient. At each iteration $b$, the weak learner is a tree with terminal nodes corresponding to phase-space partitioning—a set of non-overlapping regions $\mathcal{J}^{(b)}$ that together cover $\mathcal{X}$. This leads to
$\hat{\Delta}_{A}(\boldsymbol{x})=\sum_{b=1}^{B}\eta^{(b)}\sum_{j\in\mathcal{J}^{(b)}}\mathbbm{1}_{j}(\boldsymbol{x})\hat{\Delta}^{(b)}_{A,j}$ (57)
where the indicator function $\mathbbm{1}_{j}(\boldsymbol{x})$ equals one if $\boldsymbol{x}$ is in the phase-space region of terminal node $j$ and zero otherwise. Training iteration $b$ involves finding the partitioning $\mathcal{J}^{(b)}$ whose terminal-node predictions minimize the loss. Each terminal node prediction is based on constants $\hat{\Delta}^{(b)}_{A,j}$, which are best-fit polynomial coefficients approximating the nuisance parameter dependence in terminal node $j$. These coefficients, labeled by the multi-index $A$, are determined from the training sets $\mathcal{D}_{\mathbf{0}}$ and $\mathcal{D}^{(b)}_{\boldsymbol{\nu}}$, where we have one $\mathcal{D}^{(b)}_{\boldsymbol{\nu}}$ for each ${\boldsymbol{\nu}}\in\mathcal{V}$. We initialize with $\hat{\Delta}^{(0)}_{A,j}=0$.
To proceed from iteration $b-1$ to $b$, we remove a fraction $\eta^{(b)}$ of the previous iteration’s fit result from the training data. Since we estimate the logarithm of the DCR, the reweighting
$\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}=\left\{\exp(-\eta^{(b-1)}t^{(b-1)\ast}(\boldsymbol{x}_{i}|{\boldsymbol{\nu}}))\,w^{(b-1)}_{i},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\right\}\;\;\textrm{for all}\;\;\{w^{(b-1)},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\}\in\mathcal{D}_{{\boldsymbol{\nu}}}^{(b-1)}\;\;\textrm{for all}\;\;{\boldsymbol{\nu}}\in\mathcal{V},$ (58)
produces the corresponding $\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}$.
Algorithm 1 Boosted Parametric Tree (BPT) for learning of systematic uncertainties
@@
The nominal SM training sample $\mathcal{D}_{\mathbf{0}}$ is unchanged.
The quantity in the exponent is the best fit at iteration $b-1$,
$t^{(b-1)\ast}(\boldsymbol{x}|{\boldsymbol{\nu}})=\nu_{A}\sum_{j\in\mathcal{J}^{(b-1)}}\mathbbm{1}_{j}(\boldsymbol{x})\hat{\Delta}^{(b-1)}_{A,j},$ (59)
whose polynomial coefficients also appear on the r.h.s. of Eq. 57. To obtain $\hat{\Delta}_{A,j}^{(b)}$, we use the new training data to predict per-node cross-section values
$\sigma^{(b)}_{j,0}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}^{(b)}_{\mathbf{0}}\cap\Delta\boldsymbol{x}_{j}}w_{i}\qquad\textrm{and}\qquad\sigma^{(b)}_{j,{\boldsymbol{\nu}}}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}^{(b)}_{\boldsymbol{\nu}}\cap\Delta\boldsymbol{x}_{j}}w_{i}$ (60)
for the nominal ${\mathbf{0}}$ and each ${\boldsymbol{\nu}}\in\mathcal{V}$. The notation $\mathcal{D}^{(b)}_{\boldsymbol{\nu}}\cap\Delta\boldsymbol{x}_{j}$ indicates summing over events from $\mathcal{D}^{(b)}_{\boldsymbol{\nu}}$ that fall within the phase-space region $\Delta\boldsymbol{x}_{j}$ of terminal node $j$. These estimates, valid for ${\boldsymbol{\nu}}={\mathbf{0}}$ and ${\boldsymbol{\nu}}\in\mathcal{V}$, yield the new polynomial interpolation at iteration $b$, with coefficients